X liters of carbonmonoxide is present at STP.It is completely oxidized to co2.The volume of co2 formed is 11.207liters at STP.what is the value of X in liters?

Grade:11

1 Answers

Vikas TU
7 years ago
The balanced cheical rection is:
CO      +     1/2O2                -------------->            CO2
                                                                       moles = 11.207/22.4 = > 1.9735 
moles of CO = is also, 1.9735 from mole – 1 analysis.
Hence volume at STP for CO is:
1.9735  = X/22.4
that is same  as  = > 11.207 litre.
Hence the same Volume will be oxidized.